OpenStudy (anonymous):

partial derivatives f(x,y)= x / (x+y)^2

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can you solve for fx and fy ? i already have my own answer. i just want to compare with your answer

OpenStudy (psymon):

Oh, okay, so not much explanation needed then. Alright, lemme check.

OpenStudy (psymon):

what answers did you get?

OpenStudy (anonymous):

i got [ (x+y)^2 - 2x(x+y) ] / (x+y)^4 for fx

OpenStudy (anonymous):

0 for fy

OpenStudy (psymon):

fx is fine, fy not so much.

OpenStudy (anonymous):

hm.. i used quotient rule for fy. i thought fy was 0 since there was no 'y' on the numerator.......

OpenStudy (anonymous):

( everything is constant except for y? )

OpenStudy (psymon):

Nah, doesnt mean that at all. In fact, I wouldnt do quotient rule. If x is just a constant, I can do x(x+y)^-2 Which then I would just do chain rule = -2x(x+y)^-3 The inner derivative just becomes 1, so all I really had to do is power rule.

OpenStudy (anonymous):

so it is -2x(x+y)^3 (1) = -2x(x+y)^-3 should be answer ?

OpenStudy (psymon):

Now if I did quotient rule, just to show ya: \[\frac{ 0(x+y)^{2}-x(2)(x+y) }{ (x+y)^{4} }\]

OpenStudy (psymon):

And yep. It would come out the same way with the quotient rule, but with a slight bit more work.

OpenStudy (anonymous):

should've thought about changing the equation form.

OpenStudy (psymon):

Yeah. For me, I sometimes might literally replace the x's or y's with constants. Like, say something like "let x = 1" Then just make sure I dont lose the 1 and replace it with x in the end. Helps out a lot visually so youre not confusing yourself.

sam (.sam.):

@Psymon \[\frac{\partial \text{}}{\partial y}\left((x+y)^2\right)=2 (x+y)\]

OpenStudy (psymon):

Its coming up as an error for me right now unfortunately. Although if I did something wrong then im not surprised. I screw up more than would be wanted for someone trying to help out x_x

OpenStudy (anonymous):

thank you so much^^ im going to post up another partial derivative problem